OpenStudy (anonymous):

Okay, lets try this again... [(3^-1)(x)(y^-4)(z^5)/(7x^-4)(y^2)(z)]^-2

OpenStudy (anonymous):

Haha this looks awfully familiar. Did you have a problem with the other answer?

OpenStudy (anonymous):

I got this: \[[(x^5z^7)/(21y^6)]^-2\] so far

OpenStudy (anonymous):

Yes I did, but now I got stuck

OpenStudy (anonymous):

Mine was slightly different, the z value ended up being different from what you gave me, and I don't understand why.

OpenStudy (anonymous):

Z doesn't move. You just cancel it normally. 5-1 is 4. I didn't even pay attention to z until the very end.

OpenStudy (anonymous):

oh yea, thats right

OpenStudy (anonymous):

3^-1 right? so, where is 21 come from? I got: (x5z7)/(3y6)

OpenStudy (anonymous):

Inik the 3 moves to the bottom with the z, because it has the negative power. Then you multiply them to get 21.

OpenStudy (anonymous):

ooo... I see 7 as well. agree with 21

OpenStudy (anonymous):

Alright, l2theM, do you understand/have any more questions?

OpenStudy (anonymous):

Well, I still haven't simplified all of the way. Now that I got the above answer, how do I handle the ^-2? Do I move the top portion to the bottom, and the bottom portion to the top?

OpenStudy (anonymous):

*I did fix the z

OpenStudy (anonymous):

sorry it's messy

OpenStudy (anonymous):

Rain, how did you get the 3 on the bottom and 39 on the top?

OpenStudy (anonymous):

eh, 49 rather

OpenStudy (anonymous):

Yes. Everything gets flipped because they all become negative. Then you multiply all exponents by 2 and square the 21. You should get: \[x ^{10} z ^{8} \div 441 y ^{12}\]

OpenStudy (anonymous):

Rain, one that makes it incredibly complicated and confusing... And you also forgot to square the 3.

OpenStudy (anonymous):

I got 441y^12z^2/x^10z^10

OpenStudy (anonymous):

Yeah sorry I typed it backwards. Yours is right.

OpenStudy (anonymous):

I don't understand what you are doing with the z though. The z^5 should go to the bottom, and the z should go to the top. Why is it z^8 then?

OpenStudy (anonymous):

and also, since I have a z^2 on the top, and a z^10 on the bottom, do I simplify them to z^1/z^5?

OpenStudy (anonymous):

Because you still cancel them. 5-1 is still 4 lol Once it is cancelled you square it, and since it is an exponent you multiply them, so 4x2=8.

OpenStudy (anonymous):

Rain, I don't think that is right.

OpenStudy (anonymous):

oh it's the whole thing to the power of -2. I only did the bottom. My bad, sorry!

OpenStudy (anonymous):

You can do it either way. If you multiply first, you get the z^2/z^10 and you cancel by doing 10-2 which is still 8 lol.

OpenStudy (anonymous):

Hahaha no worries. Just remember it applies to the whole thing. I like doing things the easy way, so thats why I did it in that order.

OpenStudy (anonymous):

but why are you subtracting them if they are on opposite sides of the equation? are you changing the z^2 to z^-2 and subtracting it from the larger z^10?

OpenStudy (anonymous):

Here, let me explain another way. How would you simplify \[x ^{2} \div x\] ?

OpenStudy (anonymous):

Okay, what I did was I simplified everything in the equation and made it so all of the exponents were positive. Then I switch their locations so the ^-2 would become ^2, then I squared everything and got this: 441y^12z^2/x^10z^10

OpenStudy (anonymous):

That would be just x

OpenStudy (anonymous):

and I am sorry for the trouble, I just want to make sure I am doing this right. I have a final coming up, and I need to make a 95 or better, so really want to fully understand this.

OpenStudy (anonymous):

No worries at all. That's that this site is for! Well, you have that down, so, explain to me how you got that. And how would you do x^87/x?

OpenStudy (anonymous):

it would be x^87 with no denominator, right? I think I did the first sub-question wrong

OpenStudy (anonymous):

No you did the first one right. But what did you DO? How did you know that it was just x?

OpenStudy (anonymous):

And no, it would be x^86. so what did I do there? How did I get that answer?

OpenStudy (anonymous):

You must have just subtracted them. But I am thinking of an equivalent equation with the same bases like this: 4/2. 4 is the same as 2^2. So the equivalent equation would be 2^2/2. The answer would be 2 no matter what.

OpenStudy (anonymous):

Oh, you DID subtract the base from the numerator. x*x/x would mean 1 x goes away, which would be subtracting, right?

OpenStudy (anonymous):

Quantum, I got your answer but flipped. \[411y ^{12} \div x ^{10} z ^{8}\]What did I do wrong?

OpenStudy (anonymous):

I think you got it right this time.

OpenStudy (anonymous):

Normally I get an answer sheet with my practice tests, but my final doesn't give me one so I have no idea what the true answer is.

OpenStudy (anonymous):

I don't know how to do it by that method Rain, that looks completely foreign to me.

OpenStudy (anonymous):

Yes, that's right. You subtract them! Does that make sense now?

OpenStudy (anonymous):

Yes, it makes sense when you have x^2/x, but when you have x/x^2, that would make the problem equal x too, right?

OpenStudy (anonymous):

x/x=x?

OpenStudy (anonymous):

You subtract the bigger number by the smaller number, and whichever one was bigger, that's where it goes. Like if the x^87 was on bottom and x was on top, it would be 1/x^86. it's the same thing.

OpenStudy (anonymous):

No, x/x is one. Because both cancel each other. It's like 4/4 or 7/7.

OpenStudy (anonymous):

Ah, okay, you are using x^2 and applying a rule that the inverse would be negative in order to isolate the variables?

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

Yeah so take like, (x^2)/(x^5). what do you get?

OpenStudy (anonymous):

But if you had x/x^2, you can't subtract the top x from the bottom, because you would have a 0 in the numerator, right?

OpenStudy (anonymous):

0/x^3?

OpenStudy (anonymous):

or x^-3/0? No, that can't be right

OpenStudy (anonymous):

Yes, which would make it one. Any number^0 is one. Like y^0 is one, z^0 is one... It's always one.

OpenStudy (anonymous):

You had it right the first time, but it cancels to one. So it would be 1/x^3.

OpenStudy (anonymous):

Okay, so if you take x^2/x^5 and turned the x^2 into x^-2 and subtract it by the denominator, you assume that there is a x^0 left over, and that equals 1?

OpenStudy (anonymous):

Yes, but that could get confusing quickly. Try this. Write out all of the x's on top and bottom. Then cross out as many as you can. Then explain to me what you did and what it means.

OpenStudy (anonymous):

But essentially yes, you are correct.

OpenStudy (anonymous):

x*x/x*x*x*x*x would be 0/x^3 or 1/x^3?

OpenStudy (anonymous):

It's one. You do NOT cancel to be zero. The only way to get zero would be to subtract something. Like, you have 4 quarters in a dollar. You cancel the 4 and you got 1 dollar right? You cant just have 4 quarters and then suddenly have nothing unless someone took the money from you, ie subtracted them. Does that make sense?

OpenStudy (anonymous):

Okay yes, that makes PERFECT sense :D I think I get it now :D

OpenStudy (anonymous):

I just wanted to be sure I completely understood how you were going about this.

OpenStudy (anonymous):

Alright, make sure you got this down. Make up a problem just like the one you originally asked and simplify it. I'll tell you if you did it right.

OpenStudy (anonymous):

And again, sorry for the trouble, but I think I completely get how to go about these problems now :D I can do alot of this math, but I missed a few basic concepts back in high school, and this helps me bridge those gaps :D

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

And no worries. Haha I'm Glad to help.

OpenStudy (anonymous):

Okay, making this up... x^-5y^9z^-2/x^3y^-5z^13, going to work it out now.

OpenStudy (anonymous):

so far, x^3y^14/x^5z^15

OpenStudy (anonymous):

y^14/x^2z^15 would be the final answer

OpenStudy (anonymous):

and if I multiplied everything by ^-2, that would make it x^4z^17/y^16?

OpenStudy (anonymous):

Almost perfect. But can you tell me what is wrong with the exponent on z and y?

OpenStudy (anonymous):

for which problem?

OpenStudy (anonymous):

the first one?

OpenStudy (anonymous):

do you multiply those exponents by 2 instead of adding 2 to them?

OpenStudy (anonymous):

oh yea, x^4z^34/y^32

OpenStudy (anonymous):

thats it, got it.

OpenStudy (anonymous):

Yes, you do multiply them, you only add when you are Multiplying. Ie,( x^2)^3 = x^6 (x^2)(x^2)= x^4

OpenStudy (anonymous):

\[((3{}^{\wedge}-1)(x)(y{}^{\wedge}-4)(z{}^{\wedge}5)/(7x{}^{\wedge}-4)(y{}^{\wedge}2)(z)){}^{\wedge}-2 \] \[\left(\left(3^{-1}\right)(x)\left(y^{-4}\right)\left(z^5\right)/\left(7x^{-4}\right)\left(y^2\right)(z)\right)^{-2} \] \[\left(\left(\frac{x}{3}\right)\frac{z^5}{y^4}/\frac{7}{x^4}\left(y^2 z\right)\right)^{-2} \] \[\left(\frac{x z^5}{3 y^4}/\frac{7 y^2 z}{x^4}\right)^{-2} \] \[\left(\frac{x^5 z^4}{21 y^6}\right)^{-2} \] \[\frac{441 y^{12}}{x^{10} z^8} \]

OpenStudy (anonymous):

Lol thanks, that's been established.

OpenStudy (anonymous):

This was done with Mathmatica 8. I over looked the reference to WolframAlpha.com in a prior post above. Sorry.